) A cable operator’ quality analyst is monitoring calls from hundred customers randomly on a daily basis. She notes down instances of incorrect information being passed on to customers; the data collected over five days is given below.

Day

Monday

Tuesday

Wednesday

Thursday

Friday

Number of instances of incorrect information

5

10

23

20

15

a) Calculate the upper and lower control limits for her c-chart.

4) The specification for a Lyn's stainless steel chimney liner requires a thickness of 6.0 mm ± 0.1 mm. The firm’s quality analyst determines that the liner process operates at a mean thickness of 6.03 mm with the standard deviation of the process being 0.02 mm.

a) Calculate the upper and the lower limits for the stainless steel chimney liner.

b) Calculate the Cp and Cpk for this process.

Show the steps please!
Answer #1

a)

C-bar = Average of all number of instances

C-bar = (5+10+23+20+15)/5

C-bar = 14.6

UCL = C-bar + 3*(C-bar)^(1/2)

UCL = 14.6 + 3*(14.6)^(1/2)

UCL = 26.06

LCL = C-bar - 3*(C-bar)^(1/2)

LCL = 14.6 - 3*(14.6)^(1/2)

LCL = 3.14

4)

a)

USL = 6.0 + 0.1 = 6.1

LSL = 6.0 - 0.1 = 5.9

b)

Cp = (USL - LSL)/6*SD

Cp = (6.1-5.9)/(6*0.02)

Cp = 1.67

Cpk = Min ( (USL-Mean)/3*Sd , (Mean - LSL)/3*Sd )

Cpk = Min ( (6.1 - 6.03)/(3*0.02) , (6.03 - 5.9)/(3*0.02))

Cpk = Min (1.17 , 2.17)

Cpk = 1.17